If 2 six-sided dice are tossed, what is the probabilty that the sum of the numbers will be greater than 10?
I already know that there are three choices; 5,6 6,5 6,6
But i dont know how to set it up to solve

Answer by edjones(8007)

There are 6^2=36 possibilities
3/36=1/12 probability
.
Ed